OpenStudy (anonymous):

determine k so that the ordered pair (-3,6) is a solution of kx - 2 =9

OpenStudy (anonymous):

substitute x into the equation... k(-3)-2=9 -3k=9+2 k=-11/3

OpenStudy (anonymous):

just replace x by -3 then solve to get \[-\3x-\2\=\9\] then k is -11/3

OpenStudy (anonymous):

My answers are all whole numbers. Not fractions

OpenStudy (anonymous):

hmmm are you sure you have well written the equation??

OpenStudy (anonymous):

I screwed up it should read kx - 2y =9

OpenStudy (anonymous):

Then just substitute x=-3 and y=6 k(-3)-2(6)=9 -3k=9+12 k=21/3 k=7

OpenStudy (anonymous):

ya zats the answer